95%
The standard deviation (SD) is a measure of spread so small sd = small spread. So the above is true for any distribution, not just the Normal.
The p-value is the probability of any event or the level of significance for any statistical test.
The z-score is a transformation applied to a Random Variable with any Normal distribution to the Standard Normal distribution.
If you flip a coin 2 times how many possible outcomes are there How many combinations?
Four outcomes, three combinations.
What is the probability of getting a total of 7 when rolling two dice?
The least error prone way of doing this is to draw a 6x6 square of all the possiblities when rolling a pair of dice, where one is red and one blue, and work out the proportion of the combinations that sum to 7.
An area of a normal probability distribution represents?
The number 1. The area of any probability distribution equals 1.
A coin is tossed twice what is the probability that it will show heads both times?
The probability is 0.25.
Look at it this way--if you toss a coin twice, there are four equally-probable outcomes:
So the probability of heads twice in a row is one in four, or 25%.
the chance of tossing heads is 1/2 (50%) The chance of tossing the next heads is 1/2 (50%) 1/2 x 1/2 = 1/4 (25%)
It is 6/36 = 1/6.
There are two elements in the event space.
They are the rolls (6, 5) and (5, 6) .
What is the difference between Sub Main Distribution Board and distribution board?
main db using mcb sub main db using mccb that is deference
In a study using 9 samples, and in which the population variance is unknown, the distribution that should be used to calculate confidence intervals is
What is the chance of getting a 3 when rolling a die?
1/6 because there is only one three out of the 6 numbers on the die itself.
1/6
1 in 6 chance. Or 0.166666666666667.
With a single cube, the probability = 1/6 = 16-2/3 percent.
With a pair of cubes, the probability = 2/36 = 5-5/9 percent.
What is the probability of rolling the following numbers on a six side cube 3 4 5 or 6?
all the answers are 1/6. do you have problem with math? i can help you :)
What is the probability that N will not have Huntington disease?
0, since letters of the alphabet do not have diseases.
What is the probability that a couple with four children have 2 boys and 2 girls?
1 out of 4. Regardless of what the first two children are, there is a 50/50 chance that each of the following two kids will fulfill the remaining two conditions
Samantha Rolls died in 1996, in Vancouver, British Columbia, Canada of heroin overdose.
How many 7 of clubs in a deck of cards?
In a normal 52-card "Poker" deck, there is one "7 of Clubs" card. In fact, there's only one of every type except the Joker, and there are generally two Joker cards.
Of course, in Pinochle deck, there aren't any sevens at all.
What is the probability of rolling a number greater than 18?
The answer depends on what you are rolling: three or more ordinary dice, or fewer dice with non-standard numbers on them, or a die with some other shape.
What is the probability of 4 boys in a family with 4 children?
In a family with four children, the probability of having four boys is 1 in 16.
What is the probability of not rolling a sum of 5 with 2 dice?
It is 1 - Pr(rolling a sum of 5 with two dice)
= 1 - Pr(1+4 or 2+3 or 3+2 or 4+1) = 1 - 4/36 = 8/9